OpenStudy (anonymous):

Find all real solutions or state that there are NONE. 3e^x-9 = 6

OpenStudy (anonymous):

\[3e^x=15\] \[e^x=5\] \[x=\ln(5)\]

OpenStudy (anonymous):

Could you explain the steps please

OpenStudy (anonymous):

add 9 on both side divide both side by 3. natural log both side x = ln(5)

OpenStudy (mertsj):

ln(e^x) = ln(5) But the ln of e^x is x. That is: ln(e^x) =x. So x = ln(5)

OpenStudy (anonymous):

thank you all :D